OpenStudy (anonymous):

I'm suppose to convert r = 2/(4cos θ − 7sin θ) to Cartesian coordinates. I divided both sides by r and got 1=2/(4x+7y) but it is incorrect. What am I doing wrong?

OpenStudy (tkhunny):

Why do you believe that to be "wrong"? Perhaps if you multiplied by \(4\cos(\theta) - 7\sin(\theta)\), instead.

OpenStudy (tkhunny):

BTW, how did you get -7y rather than +7y?

OpenStudy (anonymous):

sorry that was a typo, i had -7y

OpenStudy (anonymous):

I got 4x - 7y -2 =0 as my answer now

OpenStudy (tkhunny):

Well, we do need to exclude r = 0, but other than that, does someone STILL believe it to be incorrect?

OpenStudy (anonymous):

I think it is correct now :)

OpenStudy (anonymous):

I only included the 0 because they asked to write it in an equation form.

OpenStudy (tkhunny):

No, no. That was fine all the way. My point was this... 1) In the original equation x = y = 0 is NOT allowed. 2) In the final form, that exclusion is not apparent. We have to remember it.

OpenStudy (anonymous):

ahhh I understand

OpenStudy (anonymous):

You've been helping a lot. I like how you make mee find the answer rather than telling it to me directly.

OpenStudy (tkhunny):

Note: There are other restrictions in the original equation. I'll just show you and you can ponder it when you get bored, sometime. :-) \(4\cos(x) - 7\sin(x) = \sqrt{65}\cos(x-atan(-7/4)) = 0\) Wherever that thing is zero, beings as it is in the denominator, must be excluded.

OpenStudy (anonymous):

that looks interesting, I'll ponder on it for sure.